A satellite in a force free space sweeps stationary interplanetary dust at a rate (dM/dt) = ?v. The acceleration of satellite is
    (a)    (-2av^2)/M   
    (b)    (-av^2)/M
    (c)    (-av^2)/2M    
     (d)    -av2

Grade:11

1 Answers

Navjyot Kalra
10 years ago

(b)
Thrust on the satellite,
    F=(-vdM)/dt  = -v(av)=av^2    
    Acceleration F/M=(-av^2)/M